OpenStudy (anonymous):

i will give medals if anyone can solve this question or at least attempt it!!

OpenStudy (anonymous):

Let be a continuous function defined on the interval [2, infinity[ such that f(4)=14 |f(x)| < x^3+10 and the integral from 4 to infinity f(x)*e^(-x/4) = -5 Determine the value of: the integral from 4 to infinity f'(x)*e^(-x/4) = ?

OpenStudy (anonymous):

I=\[\int\limits_{4}^{infity}( f(x)*e^(-x/4))\]=-5. I'=\[\int\limits_{4}^{infity}( f'(x)*e^(-x/4))\]=X(let) I+I'=\[\int\limits_{4}^{infity}( f(x)*e^(-x/4))\]+\[\int\limits_{4}^{infity}( f'(x)*e^(-x/4))\] -5+X=-4*([f(x)*e^x](from -1 to -infinity))=>-5+x=(56/e)=>X=(56/e)+5.....